Workshop 2 Problems Student Q Search Shoot + Review View AA. Doo Home Insert Page Layout Formulas Data P . X cut Calibri (BodWorkshop 2 Problems Student Q Search Shoot + Review View ASA. Doo Home Insert Page Layout Formulas Data P . X cut Calibri (BoCan anyone help me? Help me understand
Answer #1

Part a)

The accounting equation for different transactions is complete as below:

Donald and Friends Company
Accounting Equation
Event Assets Stockholder's Equity
Cash + Prepaid Rent = Retained Earnings
1 40,000 + 0 = 40,000
2 -20,000 + 20,000 = 0
3 0 + -18,333 (20,000*11/12) = -18,333
Total $20,000 $1,667 = $21,667

______

Part 2)

The income statement for 2018 is given as follows:

Donald and Friends Company
Income Statement
For the Year Ended December 31, 2018
Revenues 40,000
Less Expenses -18,333
Net Income $21,667

______

Part 3)

The year-end balance sheet is given as below:

Donald Duck and Friends
Balance Sheet
As of December 31, 2018
Assets
Cash 20,000
Prepaid Rent 1,667
Total Assets $21,667
Liabilities 0
Total Liabilities 0
Stockholder's Equity
Retained Earnings 21,667
Total Stockholder's Equity 21,667
Total Liabilities and Stockholder's Equity $21,667

______

Part 4)

The cash flow statement is prepared as follows:

Donald Duck and Friends
Statement of Cash Flows
As of December 31, 2018
Cash Flow from Operating Activities
Cash Receipt from Revenue 40,000
Cash Payment for Rent -20,000
Net Cash Flow from Operating Activities 20,000
Cash Flow from Investing Activities 0
Cash Flow from Financing Activities 0
Net Change in Cash 20,000
Plus: Beginning Cash Balance 0
Ending Cash Balance $20,000